# Generative models for crystalline materials

### v1

@@ Abstract
Understanding structure-property relationships in materials is fundamental in condensed matter physics and materials science. Over the past few years, machine learning (ML) has emerged as a powerful tool for advancing this understanding and accelerating materials discovery. Early ML approaches primarily focused on constructing and screening large material spaces to identify promising candidates for various applications. More recently, research efforts have increasingly shifted toward generating crystal structures using end-to-end generative models. This review analyses the current state of generative modeling for crystal structure prediction and de novo generation. It examines crystal representations, outlines the generative models used to design crystal structures, and evaluates their respective strengths and limitations. Furthermore, the review highlights experimental considerations for evaluating generated structures and provides recommendations for suitable existing software tools. Emerging topics, such as modeling disorder and defects, integration in advanced characterization, and incorporating synthetic feasibility constraints, are explored. Ultimately, this work aims to inform both experimental scientists looking to adapt suitable ML models to their specific circumstances and ML specialists seeking to understand the unique challenges related to inverse materials design and discovery.
@@ 1 Introduction
Recent reviews have broadly discussed and offered perspectives on ML-assisted inverse design in materials science [Han2025, handoko2025artificialintelligencegenerativemodels, ParkWalsh2024, wang2025crystallinematerialdiscoveryera], some of which have focused specifically on crystal structure prediction tasks [Li2024, Chen2025, Wang2025]. [debreuck2025generativeaicrystalstructures] provided a comprehensive and technical overview of the methods and models used in generative approaches for crystal structures, with discussions on methodologies and future directions, but with a primary focus on the ML models themselves [debreuck2025generativeaicrystalstructures]. [li2025materialsgenerationeraartificial] similarly emphasized the models, while also discussing datasets and evaluation metrics, and offering a brief perspective on experimental implementation, highlighting challenges such as synthesizability [li2025materialsgenerationeraartificial]. More recently, [Park2025synthesisgap] provided a perspective on the need to integrate synthesizability considerations into computational materials design [Park2025synthesisgap]. This perspective article emphasized that effective inverse design must account not only for predicted stability but also for practical pathways to experimental realization, a topic of increasing importance that will also be reviewed here.
In this work, we extend these efforts by focusing on generative models for crystalline materials, with an emphasis on recent innovations in ML approaches. We specifically focus on introducing conceptual differences of representations, datasets, and evaluation metrics, as well as discussing the practical adaptation of generative models for experimental discovery. This review should serve as a useful guide for an audience of experimental and computational materials scientists looking to find an entry point to existing generative models and how to use them in their specific applications and related materials design challenges. At the same time, this review should also inspire the materials informatics and computer science communities, looking for an experimental perspective on currently open challenges and considerations to take into account in the next generation of generative models for materials.
@@ 3 Databases
The effectiveness of ML methods strongly depends on the quality of the underlying data, as high-quality data is crucial for ensuring model reliability and generalization. The development of comprehensive and accurate crystal structure databases is thus of paramount importance in computational materials science, motivating the community efforts to invest in curated databases. On the one hand, the experimental materials science community has been gathering measurements and characterization results into open-access databases. On the other hand, the scarcity of experimental data has motivated the generation of large-scale simulated datasets using DFT, which complement and expand existing resources. While these databases are often categorized as experimental or DFT-calculated, many are hybrid in practice. Additional efforts have also led to the creation of application-specific databases tailored to particular materials. In this section, we present existing publicly available crystal structure databases, also summarized in Table 1.
@@
Materials Project (Full) [Jain2013] Mixed (some verified) >700,000 total Molecules and inorganic crystals
@@
Table 1: Summary of major publicly available materials databases from the literature, highlighting key characteristics including database size, maximum number of atoms per unit cell, and material types covered. Experimental validation is indicated where applicable — while some databases are purely computational, others combine theoretical and experimentally verified structures to varying degrees.
@@ 3.3 Application-specific databases
For catalysts, the OC20 dataset, consisting of 1,281,040 relaxations of DFT relaxations (264,890,000 single-point evaluations) across a wide swath of materials, surfaces, and adsorbates (nitrogen, carbon, and oxygen chemistries) [chanussot2021open]. The Northeast Materials Database (NEMAD) is a database that consists of 67,573 magnetic materials collected via LLM-based automated scientific data extraction from experimental literature [zhang2024gptarticleextractor, itani2024northeast]. Supercon focuses on high-$T_{c}$ oxide superconductors, metal-alloy superconductors, and organic superconductors, with data extracted from academic papers and other sources [materials-a], complemented by 3DSC, a dataset of superconductors based on Supercon but including matched crystal structures from the ICSD and Materials Project [sommer20233dsc].
@@ 4.1 Design by combining property prediction with high-throughput screening and search
Taken together, these DFT-based, ML-assisted, and heuristic search methods form the foundation of modern CSP and de novo generation workflows. Figure 3 provides an overview of this landscape, summarizing how property prediction, search-based generation, and end-to-end generative models are connected, along with literature spotlights. These approaches illustrate a transition from purely physics-based prediction to data-augmented search and screening. These frameworks generate large volumes of plausible structures, evaluate them with ML and DFT, and refine candidates toward realistic, stable materials. This layered approach—generate, predict, validate—sets the stage for the emergence of end-to-end generative models (Figure 4), which now aim to invert the process entirely: to learn how to directly propose structures with desired properties, a topic explored in the following sections.
@@ 5.2 Usability: software and code availability
Access to open-source, ready-to-use software is crucial for researchers looking to apply generative models in practice. To provide a practical starting point, we grouped a selection of implementations that we consider to be scientifically relevant and representative of recent advances in the field. This selection is not complete and balances multiple criteria beyond usual benchmark metrics, including code documentation and usability. Table 2 summarizes a selection of recent generative models for materials, including their associated datasets, code availability, checkpoints, and conditioning options.
@@
MatterGen [zeni2025GenerativeModelInorganic] Diffusion
Alex-MP-20,
MP-20
GitHub HuggingFace
Composition, Space group,
Band gap, Bulk modulus,
Magnetic density, HHI score,
Energy above hull
DiffCSP [jiao2024CrystalStructurePrediction] Diffusion
Carbon, Perov,
MP-20
GitHub Google Drive Composition, Formation Energy
DiffCSP++ [jiao2024SpaceGroupConstrained] Diffusion
Carbon, Perov,
MP-20, MPTS-52
GitHub Google Drive Space group constrained. Composition
CrysBFN [wu2025PeriodicBayesianFlow] BFN
Carbon, Perov,
MP-20, MPTS-52
GitHub Google Drive Composition
CrystalFormer [cao2024SpaceGroupInformed] Transformer
Alex-20,
MP-20
GitHub HuggingFace
Space group constrained.
Dielectric FoM, Energy above hull
(with reinforcement fine-tuning)
Table 2: Overview of selected repositories of generative models for materials.
@@ 5.3 Speed and computational cost
Inference speed is a critical consideration in the practical deployment of generative models for crystal structure generation. A variation in inference time exists among methods, largely due to differences in architectural complexity, data representation, and sampling requirements [ruple2025symmetryawarebayesianflownetworks]. Exact inference runtimes depend heavily on hardware configurations, so comparisons between different methods are limited to estimates at the level of orders of magnitude. Diffusion-based generative models generally exhibit the longest inference times due to the requirement of thousands of iterative denoising steps during sampling [jiao2024CrystalStructurePrediction]. Fine-tuned LLMs also operate at a high computational cost, with inference speeds varying over multiple orders of magnitude depending on model size. This variation reflects a tradeoff between computational efficiency and the quality of the generated structures [gruver2024FineTunedLanguageModels], where the largest LLMs approach inference times comparable to those of diffusion models. Flow-based models operate at a similar speed to the smallest fine-tuned LLMs, as they generally require fewer sampling steps than diffusion models for generation [millerFlowMMGeneratingMaterials2024a]. Smaller transformer-based architectures employing symmetry-aware tokenization schemes achieve inference speeds approximately an order of magnitude faster than the smallest fine-tuned LLMs [cao2024SpaceGroupInformed]. BFNs improve upon the number of required sampling steps, reducing inference time by yet another order of magnitude compared to the transformer models [ruple2025symmetryawarebayesianflownetworks], bringing the inference time to generate a crystal structure down to a few milliseconds.
@@ 5.5 Post generation selection workflows
The minimum assessment for synthesizability is phase stability, agnostic of experimental method and condition. In other words, how likely is it that the generated structure is real? In Figure 6, we show a filtration schema for phase stability assembled from collections of target prediction and synthesizability literature. The criteria of (i) energetic stability, (ii) chemical heuristics, and (iii) structural stability are organized by approximate resource requirements, e.g., least to most computationally and time-intensive. As previously discussed in Section 5.4, many screening workflows search initially for favorable (thermodynamic) energetics, often evaluated by a negative formation energy ($\Delta H_{f}$), energy ($E_{\mathrm{hull}}$) with respect to the convex (composition-energy) hull, or appearance as the most stable phase on a phase diagram e.g. (composition-temperature or chemical potential diagrams) [bartel2022ReviewComputationalApproaches, Ong2008_ConvexHull]. The former parameters are easy to integrate into automatic workflows leveraging computational materials databases of bulk crystalline inorganic solids, such as the Materials Project and the Open Quantum Materials Database [Jain2013, Horton2025_MaterProj_2, Saal2013_OQMD_1, kirklin2015open]. Nearly all known bulk inorganic solids exhibit negative formation energies sitting within 94 meV/atom of the convex hull, making $E_{\mathrm{hull}}$ a tool useful for quickly identifying materials unlikely to form [Sun2016_Metastability_1, Aykol2018_Metastability_2]. However, the threshold for metastability has been shown to be chemistry and chemical system dependent and thus prevents it from being a one-size-fits-all metric [Aykol2018_Metastability_2]. Screening for chemical heuristics in proposed structures ensures that critical chemical laws known to constrain ion oxidation state, polyhedra geometries, bonding orders, and coordination numbers are followed. Historically, the evaluation of chemical rules of thumb can be challenging to enforce due to the lack of formal, rigid rules governing bonding and structure for translationally invariant bulk, crystalline solids. However, a number of studies have successfully employed specific heuristic metrics to known and candidate structures, including the Pauling rules, the Crystal Orbital Bond Index, structure-type tolerance factors, oxidation state predictors, and the mno counting rules for intermetallic bonding [Mller2021_COBI, George2020_PaulingRules, Simonson2010_BondingScheme, Filip2018_GoldschmidtTolFac, Fu2023_OSP_1, Thway2024_OSP_2]. Structural thermal dynamics and stability are important to understand proposed phase feasibility, and, in particular, towards assessing relative stability to a ground state atomic configuration. Structural stability can be measured through Monte Carlo (simulated annealing), molecular dynamics, and phonon calculations [Jansen2018_moreModernReview, Jansen2002_StructureEnergyLandscape, Malyi2019_EngPhDynStability, Qian2021_PhononRev, Ozoli1999_PhononScreening], though computational and time resources are likely to be higher than for other screening metrics. It is important to note that many known materials may exhibit some lattice instability, such as a high-temperature metastable polymorph, and as such, alongside high resource costs, structural dynamics screening may not be included in high-throughput structure prediction studies. Passing through one or more of these screening steps provides researchers with some confidence in the phase stability and physicality of a generated structure.
@@ 5.6 Emerging topics
While most of the previous generative approaches focus on ideal crystalline materials, real solids can contain defects and disorder due to, e.g., vacancies, substitutions, and dislocations. As these features can affect the final properties of the crystals, accounting for them is essential to bridge the gap between theoretical predictions and experimentally realizable materials. ML and generative models for engineering defects and incorporating disorder represent an important design opportunity for broad classes of functional materials, such as 2D materials, ferroelectrics, and small-gap semiconductors, wherein computational crystal structure generation was previously tedious and incomplete. Recent studies have begun addressing this challenge by developing models and frameworks capable of explicitly representing and learning from non-ideal structures. [Yang2025_DefectEng_1] introduced DefiNet, a defect-informed EGNN designed to capture the local geometries and interactions in point-defect structures [Yang2025_DefectEng_1]. [Frey2020_DefectEng_2] applied deep transfer learning to predict defect properties in 2D materials, identifying numerous dopant–defect combinations with potential for quantum emission and neuromorphic computing [Frey2020_DefectEng_2]. Mosquera-Lois et al. used a machine-learning surrogate model to accelerate defect structure relaxation, successfully predicting stable reconstructions for unseen materials and significantly reducing DFT [MosqueraLois2024_DefectEng_3]. Beyond point defects, [Petersen2025_DefectEng_4] developed Dis-GEN, the first generative model capable of producing symmetry-consistent disordered crystal structures with partial occupancies and vacancies, trained directly on ICSD data [Petersen2025_DefectEng_4]. Finally, [Jakob2025_DefectEng_5] introduced ML classifiers to predict the likelihood of crystallographic disorder in inorganic compounds, providing a path toward disorder-aware materials discovery workflows that better mirror experimental realities [Jakob2025_DefectEng_5]. Together, these developments mark an important shift from modeling perfect crystals toward embracing the inherent complexity of real materials by incorporating defects and disorder.

### v2

# arXiv:2511.22652v2[cond-mat.mtrl-sci] 02 Mar 2026

@@ Abstract
Understanding structure-property relationships in materials is fundamental in condensed matter physics and materials science. Over the past few years, machine learning (ML) has emerged as a powerful tool for advancing this understanding and accelerating materials discovery. Early ML approaches primarily focused on constructing and screening large material spaces to identify promising candidates for various applications. More recently, research efforts have increasingly shifted toward generating crystal structures using end-to-end generative models. This review analyzes the current state of generative modeling for crystal structure prediction and de novo generation. It examines crystal representations, outlines the generative models used to design crystal structures, and evaluates their respective strengths and limitations. Furthermore, the review highlights experimental considerations for evaluating generated structures and provides recommendations for suitable existing software tools. Emerging topics, such as modeling disorder and defects, integration in advanced characterization, incorporating synthetic feasibility constraints, and model explainability are explored. Ultimately, this work aims to inform both experimental scientists looking to adapt suitable ML models to their specific circumstances and ML specialists seeking to understand the unique challenges related to inverse materials design and discovery.
@@ 1 Introduction
Recent reviews have broadly discussed and offered perspectives on ML-assisted inverse design in materials science [Han2025, handoko2025artificialintelligencegenerativemodels, ParkWalsh2024, wang2025crystallinematerialdiscoveryera, Cheng2026], some of which have focused specifically on crystal structure prediction tasks [Li2024, Chen2025, Wang2025]. [debreuck2025generativeaicrystalstructures] provided a comprehensive and technical overview of the methods and models used in generative approaches for crystal structures, with discussions on methodologies and future directions, but with a primary focus on the ML models themselves [debreuck2025generativeaicrystalstructures]. [li2025materialsgenerationeraartificial] similarly emphasized the models, while also discussing datasets and evaluation metrics, and offering a brief perspective on experimental implementation, highlighting challenges such as synthesizability [li2025materialsgenerationeraartificial]. More recently, [Park2025synthesisgap] provided a perspective on the need to integrate synthesizability considerations into computational materials design [Park2025synthesisgap]. This perspective article emphasized that effective inverse design must account not only for predicted stability but also for practical pathways to experimental realization, a topic of increasing importance that will also be reviewed here.
In this work, we extend these efforts in multiple aspects by focusing on generative models for crystalline materials with (i) an emphasis on recent innovations in ML approaches in the last years, i.e., models based on diffusion, flows, and transformers; (ii) a specific focus on introducing conceptual differences of representations, datasets, and evaluation metrics, also making these distinctions accessible to non-experts and new researchers in this field; and (iii) to enhance practical relevance beyond previous review papers, a discussion of the practical adaptation of generative models for experimental discovery, as well as the current limitations and emerging topics, e.g., defects and disorder, characterization workflow integration, synthetic feasibility, and model interpretability.
Combining these three aspects, this review should serve as a useful guide for an audience of experimental and computational materials scientists looking to find an entry point to existing generative models and how to use them in their specific applications and related materials design challenges. At the same time, this review should also inspire the materials informatics and computer science communities, looking for an experimental perspective on currently open challenges and considerations to take into account in the next generation of generative models for materials.
@@ 3 Databases
The effectiveness of ML methods strongly depends on the quality of the underlying data, as high-quality data is crucial for ensuring model reliability and generalization. The development of comprehensive and accurate crystal structure databases is thus of paramount importance in computational materials science, motivating the community efforts to invest in curated databases. On the one hand, the experimental materials science community has been gathering measurements and characterization results into open-access databases. On the other hand, the scarcity of experimental data has motivated the generation of large-scale simulated datasets using DFT, which complement and expand existing resources. While these databases are often categorized as experimental or DFT-calculated, many are hybrid in practice. Additional efforts have also led to the creation of application-specific databases tailored to particular materials. It is important to note that existing crystal structure databases are inherently imbalanced. Certain chemical elements, space groups, and structure types are strongly overrepresented due to historical research focus, experimental accessibility, and computational screening strategies. These representation biases can influence generative models in the absence of additional constraints. However, the quantitative impact of these biases remains difficult to assess given the lack of alternative, unbiased reference datasets. In this section, we present existing publicly available crystal structure databases, also summarized in Table 1.
@@
Materials Project (Full) [Jain2013] Mixed (some verified) >700,000 total, 200,000 crystals Molecules and inorganic crystals
@@
Table 1: Summary of major publicly available materials databases from the literature, highlighting key characteristics including database size and material types covered. Experimental validation is indicated where applicable — while some databases are purely computational, others combine theoretical and experimentally verified structures to varying degrees.
@@ 3.3 Application-specific databases
For catalysts, the OC20 dataset consists of 1,281,040 DFT relaxations (264,890,000 single-point evaluations) across a wide swath of materials, surfaces, and adsorbates (nitrogen, carbon, and oxygen chemistries) [chanussot2021open]. The Northeast Materials Database (NEMAD) is a database that consists of 67,573 magnetic materials collected via LLM-based automated scientific data extraction from experimental literature [zhang2024gptarticleextractor, itani2024northeast]. Supercon focuses on high-$T_{c}$ oxide superconductors, metal-alloy superconductors, and organic superconductors, with data extracted from academic papers and other sources [materials-a], complemented by 3DSC, a dataset of superconductors based on Supercon but including matched crystal structures from the ICSD and Materials Project [sommer20233dsc].
@@ 4.1 Design by combining property prediction with high-throughput screening and search
Taken together, these DFT-based, ML-assisted, and heuristic search methods form the foundation of modern CSP and de novo generation workflows. Figure 3 provides an overview of this landscape, summarizing how property prediction, search-based generation, and end-to-end generative models are connected, along with literature spotlights. These approaches illustrate a transition from purely physics-based prediction to data-augmented search and screening. These frameworks generate large volumes of plausible structures, evaluate them with ML and DFT, and refine candidates toward realistic, stable materials. This layered approach—generate, predict, validate—sets the stage for the emergence of end-to-end generative models (Figure 4), which now aim to invert the process entirely: to learn how to directly propose structures with desired properties, a topic explored in the following sections. In Table 2, we provide a first, concise overview of major end-to-end generative model families for crystal structure design, highlighting key trade-offs in generation quality, inference speed, training stability, and likelihood estimation.
Model Generation quality Inference speed Training stability Likelihood estimation
Variational autoencoders Medium Fast: One-shot generation Often hard to tune Only ELBO
Generative adversarial networks Medium Fast: One-shot generation Often prone to minimax oscillations Not available
Diffusion models Competing for SOTA Typically slow: Many iterative steps Usually stable: Score matching objectives; can be sensitive to noise scheduling Possible but very expensive
Flow matching Medium: Typically fewer steps than diffusion Usually stable: Flow-matching objectives Possible but very expensive
Bayesian flow networks Medium to high: Typically fewer steps than diffusion Usually stable: Bayesian update framework Not reported in literature
Large language models Typically slow to medium: Token-by-token generation Usually stable: Standard cross-entropy objective Only auto-regressive likelihoods
Table 2: Qualitative comparison of model families for crystal structure generation. The entries provide approximate assessments intended to guide intuition rather than quantitative ranking. Actual performance depends strongly on representation, sampling steps, model and batch size, hardware, implementation details, and related experimental choices.
@@ 5.2 Usability: software and code availability
Access to open-source, ready-to-use software is crucial for researchers looking to apply generative models in practice. To provide a practical starting point, we grouped a selection of implementations that we consider to be scientifically relevant and representative of recent advances in the field. This selection is not complete and balances multiple criteria beyond usual benchmark metrics, including code documentation and usability. Table 3 summarizes this selection of recent generative models for materials, including their associated datasets, code availability, checkpoints, and conditioning options.
@@
MatterGen [zeni2025GenerativeModelInorganic] Diffusion Alex-MP-20, MP-20 GitHub HuggingFace Composition, Space group, Band gap, Bulk modulus, Magnetic density, HHI score, Energy above hull
DiffCSP [jiao2024CrystalStructurePrediction] Diffusion Carbon, Perov, MP-20 GitHub Google Drive Composition, Formation Energy
DiffCSP++ [jiao2024SpaceGroupConstrained] Diffusion Carbon, Perov, MP-20, MPTS-52 GitHub Google Drive Space group constrained. Composition
CrysBFN [wu2025PeriodicBayesianFlow] BFN Carbon, Perov, MP-20, MPTS-52 GitHub Google Drive Composition
CrystalFormer [cao2024SpaceGroupInformed] Transformer Alex-20, MP-20 GitHub HuggingFace Space group constrained. Dielectric FoM, Energy above hull (with reinforcement fine-tuning)
Table 3: Overview of selected repositories of generative models for materials. Repositories were selected based on competitive (near state-of-the-art) performance and the availability of well-documented, reproducible code and pretrained checkpoints.
@@ 5.3 Speed and computational cost
Inference speed is a critical consideration in the practical deployment of generative models for crystal structure generation. A variation in inference time exists among methods, largely due to differences in architectural complexity, data representation, and sampling requirements [ruple2025symmetryawarebayesianflownetworks]. Exact inference runtimes depend heavily on hardware configurations, so comparisons between different methods are limited to estimates at the level of orders of magnitude. For iterative models, the number of sampling steps is the primary determinant of runtime, since each step involves one or two computationally expensive neural network evaluations. In many reported settings, diffusion-based generative models exhibit long inference times because achieving state-of-the-art performance often involves hundreds to thousands of iterative denoising steps during sampling [jiao2024CrystalStructurePrediction, levy2025SymmCDSymmetryPreservingCrystal]. Fine-tuned LLMs also operate at a high computational cost, with inference speeds varying over multiple orders of magnitude depending on model size. This variation reflects a tradeoff between computational efficiency and the quality of the generated structures [gruver2024FineTunedLanguageModels], where the largest LLMs approach inference times comparable to those of diffusion models. Flow-based models operate at a similar speed to the smallest fine-tuned LLMs, as they generally require fewer sampling steps than diffusion models for generation, typically on the order of a few hundred steps [millerFlowMMGeneratingMaterials2024a]. Smaller transformer-based architectures employing symmetry-aware tokenization schemes achieve inference speeds approximately an order of magnitude faster than the smallest fine-tuned LLMs [cao2024SpaceGroupInformed]. BFNs improve upon the number of required sampling steps, typically using fewer than 100 steps, reducing inference time by yet another order of magnitude compared to the transformer models. With the experimental setup reported for SymmBFN [ruple2025symmetryawarebayesianflownetworks], using 100 sampling steps corresponds to generating a crystal structure in a few milliseconds on NVIDIA A100 GPUs.
@@ 5.5 Post generation selection workflows
The minimum assessment for synthesizability is phase stability, agnostic of experimental method and condition. In other words, how likely is it that the generated structure is real? In Figure 6, we show a filtration schema for phase stability assembled from collections of target prediction and synthesizability literature. The criteria of (i) energetic stability, (ii) chemical heuristics, and (iii) structural stability are organized by approximate resource requirements, e.g., least to most computationally and time-intensive. As previously discussed in Section 5.4, many screening workflows search initially for favorable (thermodynamic) energetics, often evaluated by a negative formation energy ($\Delta H_{f}$), energy ($E_{\mathrm{hull}}$) with respect to the convex (composition-energy) hull, or appearance as the most stable phase on a phase diagram e.g. (composition-temperature or chemical potential diagrams) [bartel2022ReviewComputationalApproaches, Ong2008_ConvexHull]. The former parameters are easy to integrate into automatic workflows leveraging computational materials databases of bulk crystalline inorganic solids, such as the Materials Project and the Open Quantum Materials Database [Jain2013, Horton2025_MaterProj_2, Saal2013_OQMD_1, kirklin2015open]. Nearly all observed bulk inorganic solids have negative formation energies and are either on the hull or are within $\approx$ 100meV/atom of the convex hull, making $E_{\mathrm{hull}}$ a useful tool [Sun2016_Metastability_1, Aykol2018_Metastability_2]. However, the threshold for metastability has been shown to be chemistry and chemical system dependent and thus prevents it from being a one-size-fits-all metric [Aykol2018_Metastability_2]. Screening for chemical heuristics in proposed structures ensures that critical chemical laws known to constrain ion oxidation state, polyhedra geometries, bonding orders, and coordination numbers are followed. Historically, the evaluation of chemical rules of thumb can be challenging to enforce due to the lack of formal, rigid rules governing bonding and structure for translationally invariant bulk, crystalline solids. However, a number of studies have successfully employed specific heuristic metrics to known and candidate structures, including the Pauling rules, the Crystal Orbital Bond Index, structure-type tolerance factors, oxidation state predictors, and the mno counting rules for intermetallic bonding [Mller2021_COBI, George2020_PaulingRules, Simonson2010_BondingScheme, Filip2018_GoldschmidtTolFac, Fu2023_OSP_1, Thway2024_OSP_2]. Structural thermal dynamics and stability are important to understand proposed phase feasibility, and, in particular, towards assessing relative stability to a ground state atomic configuration. Structural stability can be measured through Monte Carlo (simulated annealing), molecular dynamics, and phonon calculations [Jansen2018_moreModernReview, Jansen2002_StructureEnergyLandscape, Malyi2019_EngPhDynStability, Qian2021_PhononRev, Ozoli1999_PhononScreening], though computational and time resources are likely to be higher than for other screening metrics. It is important to note that many high-temperature metastable polymorphs may exhibit some lattice instability as the temperature is lowered and are therefore challenging to evaluate with high-throughput 0 K calculations, even when phonons are considered.
@@
The lack of a theory of synthesizability is particularly challenging for generative structure models, as synthesis feasibility cannot be included as a constraint during the generation, but can at best be evaluated afterwards through explicit calculations of convex hulls and phase diagrams. In principle, requirements for the topology of the convex hull near the generated compound could be included during structure generation, though this is complex due to the non-local nature of hull energy, depending on energies at many other compositions, and has therefore never been demonstrated as a feasible constraint in generative structure generation.
@@ 5.6 Emerging topics
While most of the previous generative approaches focus on ideal crystalline materials, real solids can contain defects and disorder due to, e.g., vacancies, substitutions, and dislocations. As these features can affect the final properties of the crystals, accounting for them is essential to bridge the gap between theoretical predictions and experimentally realizable materials. ML and generative models for engineering defects and incorporating disorder represent an important design opportunity for broad classes of functional materials, such as 2D materials, ferroelectrics, and small-gap semiconductors, wherein computational crystal structure generation was previously tedious and incomplete. Recent studies have begun addressing this challenge by developing models and frameworks capable of explicitly representing and learning from non-ideal structures. It is worth noting that, to date, most approaches in the literature addressing defects and disorder rely on conventional ML models with the ability to complement generative frameworks through screening, prediction, or post-processing. Only a limited number of works, to the best of our knowledge, directly incorporate defects or disorder within the structure generation process itself.
[Yang2025_DefectEng_1] introduced DefiNet, a defect-informed EGNN designed to capture the local geometries and interactions in point-defect structures [Yang2025_DefectEng_1]. A key distinction of this approach relative to conventional GNNs is the use of an explicit defect representation rather than relying on defects being implicitly inferred from atomic geometry. Each node in the graph carries additional markers that encode whether the atom corresponds to a pristine lattice site, a substitutional defect, or a vacancy, making defect identity directly available during message passing. This defect-aware formulation allows the network to learn how different defect types locally perturb atomic interactions and structural relaxations. [Frey2020_DefectEng_2] applied deep transfer learning to predict defect properties in 2D materials, identifying numerous dopant–defect combinations with potential for quantum emission and neuromorphic computing [Frey2020_DefectEng_2]. Defect effects are incorporated through engineered feature vectors that explicitly encode local chemical and structural deviations from the pristine lattice, enabling a classifier to predict defect occurrence and a regressor to estimate defect formation energies. Mosquera-Lois et al. used a machine-learning surrogate model to accelerate defect structure relaxation, successfully predicting stable reconstructions for unseen materials and significantly reducing DFT [MosqueraLois2024_DefectEng_3]. This framework focuses on post-generation structural refinement around defects, rather than producing full atomic configurations from scratch. Beyond point defects, [Jakob2025_DefectEng_5] introduced ML classifiers to predict the likelihood of crystallographic disorder in inorganic compounds, providing a path toward disorder-aware materials discovery workflows that better mirror experimental realities [Jakob2025_DefectEng_5]. [Petersen2025_DefectEng_4] developed Dis-GEN, the first generative model capable of producing symmetry-consistent disordered crystal structures with partial occupancies and vacancies, trained directly on ICSD data [Petersen2025_DefectEng_4]. Dis-GEN is a VAE-based framework that explicitly represents occupational disorder and vacancies at the level of crystallographic sites. Disorder is encoded at each Wyckoff position together with a binary indicator specifying whether a site is ordered or disordered. This enables the generation of configurationally disordered structures that more faithfully reflect experimental crystallography. Together, these developments mark an important shift from modeling perfect crystals toward embracing the inherent complexity of real materials by incorporating defects and disorder. However, the widespread adoption of defect- and disorder-aware generative models remains constrained by the scarcity of data on defects and disordered materials, making this a major bottleneck for inverse design workflows.
@@
Model interpretability and explainability
In recent years, the question of model interpretability has emerged as a central topic in machine learning research [longoExplainableArtificialIntelligence2024a]. The ever-increasing complexity of modern models and their resulting incomprehensible black-box character have raised concerns about user trust, technical pitfalls such as artifacts and hallucinations, and societal concerns about model bias and fairness. Research in explainable AI aims to develop methods to mitigate these risks by increasing the transparency of these black box models. While not as well explored as for predictive models, some research exists on the interpretability of generative models [schneiderExplainableGenerativeAI2024], centered around meaningful latent space decompositions [moranInterpretableDeepGenerative2025], identification of particularly influential training samples [nguyenRegionLevelDataAttribution2025], and model parameters [nguyenUnveilingConceptAttribution2024]. While these topics are being actively explored in the greater context of generative modeling, the aspect of interpretability is still notably underrepresented in crystal structure generation, aside from some isolated discussion of latent space decomposition [wangInverseDesignCatalytic2025, court3DInorganicCrystal2020], demonstrating an important direction for future developments of the field.